1/4

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Der Wer- wolf Le loup-garou Varulven The Bans- hee (An Approach) The Were- wolf

übersetzt von R. Platteau übersetzt von Sven Collberg übersetzt von Max Knight übersetzt von Alexander Gross

Ein Werwolf eines Nachts entwich

von Weib und Kind und sich begab

an eines Dorfschullehrers Grab

und bat ihn: �Bitte, beuge mich!“ Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école

Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Der Dorfschulmeister stieg hinauf

auf seines Blechschilds Messingknauf

und sprach zum Wolf, der seine Pfoten

geduldig kreuzte vor dem Toten:

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école

Et lui dit, “ sur moi dis tout, c’est ton rôle ! ” mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp

med svar till den, som satt på pass med tåligt korslagd tass på tass.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

…  …  …  …  …

2/4

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Genitiv sodann,

dem Werwolf, Dativ, wie man’s nennt,

den Werwolf, - damit hat’s ein End!“ “ - On peut à ton sujet puisque nous y sommes

Se demander encor bien des choses en somme.

Loup-garoù ? mais

avant Il faut se dire : loupgarquand ?

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case– and that is all.”

“Well, ‘Werewolf’ is your plural past,

While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense,

And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

Dem Werwolf schmeichelten die Fälle

er rollte seine

Augenbälle. �Indessen“, bat

er, �füge doch

zur Einzahl auch die Mehrzahl noch!“ Loup-garcomment ?

et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au loup bien absconses, Car en fait ce qu’il voulait c’était les réponses.

Varulven myste glad åt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied.

The scholar paused–again he tried:

Der Dorfschullehrer aber musste

gestehn, dass er von ihr nichts wusste.

Zwar Wölfe gab’s in großer Schar,

doch �Wer“ gab’s nur im Sinular.

“ - Point de réponses, dit le puits de savoir,

Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes.

“ - Oh!, même les baisers de ma douce moitié Bekänna

måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong!

Such words are wounds beyond all suture– I’m sorry, but you have no future.”

…  …  …  …  …

3/4

Der Wolf erhob sich tränenblind –

er hatte ja doch Weib und Kind!

Doch da er keine Gelehrter eben,

so schied er dankend und ergeben.

Ne pourrons plus me faire me sentir tout entier. Rester un loup-garou est un sort qui m’alarme. ” Mais comme le loup était bien poli Il remercia et puis re-

partit.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

The banshee, rising clammily,

wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

The Werewolf knew better– his sons still slept

At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

…  …  …  …  …

4/4

Verfasser dieses deutschen Gedichtes ist Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Französische durch R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu  ĉi  franci-
go  troviĝas  en

http: // bertilow. com/ literaturo/ lupfantomo. html .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Schwedische durch Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Englische durch Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_ poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Englische durch Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu  ĉi  angli-

go troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .